OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1 fourthx + 2.

OpenStudy (anonymous):

I will give 10 medals

OpenStudy (michele_laino):

please you have to apply this formula: \[y-y _{0}=m(x-x _{0})\] where (x_0,y_0) is your point and m is your slope, namely 1/4

OpenStudy (anonymous):

these are the answer choices y = 4x - 1 y = 4x - 3 y = 1 fourthx - 3 y = 1 fourthx - 1

OpenStudy (anonymous):

which one do u think it is

OpenStudy (zale101):

The slope is 1/4 because you are looking for the parallel line of the equation y=1/4+2. Parallel lines have the same slope.

OpenStudy (zale101):

y=1/4x+2*

OpenStudy (michele_laino):

that's right! @Zale101

OpenStudy (zale101):

Based off of what michele and i said, what you think the answer is? @miamelii

OpenStudy (anonymous):

I have no idea haha

OpenStudy (zale101):

You're not trying :\

OpenStudy (anonymous):

I mean I dont understand it, I think its C

OpenStudy (zale101):

Apply the point slope formula given by michele. Plug in the given point (-4,-2) and the slope 1/4 for m. Have you tried that?

OpenStudy (anonymous):

yes so its B

Nnesha (nnesha):

like they already said parallel mean SAME slope do you know what is slope of this line y = 1 fourthx + 2 ???

OpenStudy (anonymous):

what?

OpenStudy (alexandervonhumboldt2):

the slope are the same of parallel lines so the slope of the line is 1/4. plug in -4, -2 -2=1/4*-4+b b=-1 y=1/4x-1

Nnesha (nnesha):

y = mx + b slope intercept form equation where m is slope which is always with x and b is y-intercept equation you have is already in slope form so what is m in your equation what number is with x ???

OpenStudy (anonymous):

1/4

OpenStudy (alexandervonhumboldt2):

ys

Nnesha (nnesha):

yes right now you can cancel out two answer choices

OpenStudy (anonymous):

thankyou @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

k

OpenStudy (anonymous):

wait so the anser is D

Nnesha (nnesha):

which two answer choices you can cancel ???

OpenStudy (anonymous):

A/B

Nnesha (nnesha):

yes right

OpenStudy (alexandervonhumboldt2):

yes d

Nnesha (nnesha):

now use same slope intercept form equation they gave you x and y values now replace x and y by -4 and -2 then solve for b